OpenStudy (anonymous):

solve by graphing x+y=6 x-y=4

OpenStudy (campbell_st):

well use a table of values for each line x + y = 6... or y = 6 - x x: 0 : 1: 2 --------- y: 6 : 5: 4 plot the points (0,6), (1, 5), (2, 4) on the number plane same process for your other line x - y = 4 or y = x - 4 x: 0 : 1: 2 --------- y: -4: -3: -2 plot the ordered pairs the point of intersection is the solution

OpenStudy (luigi0210):

what kind of graph do you want?

OpenStudy (luigi0210):

hand drawn or a colorful graph?

OpenStudy (anonymous):